OpenStudy (anonymous):

A geometric sequence is 3/4, 9, 108, 1296, 15,552, 186,624, ... . Which is the general term of the sequence? A)tn = 3/32(12)n - 1, where n ∈N and n ≥1 B)tn = 3/16(12)n - 1, where n ∈N and n ≥1 C)tn = 3/8(12)n - 1, where n ∈N and n ≥1 D)tn = 3/4(12)n - 1, where n ∈N and n ≥1

OpenStudy (anonymous):

a = 3/4 ar = 9 r = \[9\div3/4\] = 12 tn = ar^(n-1) general formula tn = (3/4) 12^(n-1) Answer is D

OpenStudy (anonymous):

D